y=1/2x +4
2y+2x=2

Answer by elima(1433)

y=1/2x +4
2y+2x=2
Solve this by substitution;
Substitute y in eq 2 for y=1/2x+4
2(1/2x+4)+2x=2
x+8+2x=2
3x=-6
x=-2
Now solve the eq. with the answer you got for x;
2y+2(-2)=2
2y-4=2
2y=6
y=3
So now you have;
x=-2
y=3
Hope you understand
=)